A culture of bacteria has 20 cells in it the number of cells doubles every 30 minutes.
A)use the geometric general term formula to find the number of cells after 8 hours

Answer by stanbon(75887)

A culture of bacteria has 20 cells in it the number of cells doubles every 30 minutes.
A)use the geometric general term formula to find the number of cells after 8 hours
====
n = ab^x
------
n = 20*(2)^(h/(1/2))
------
n = 20*2^(8/(1/2))
------
n = 20*2^(16)
------
n = 20*2^16 cells = 1310720 cells
------------
Cheers,
Stan H.
---------------